# Constructing bounded orbits of special types on homogeneous spaces

### v1

@@ Abstract.
Let $X=G/\Gamma$ be a quotient of a real Lie group by a non-uniform lattice. Consider a one-parameter subgroup $F$ of $G$ that is $\operatorname{Ad}$-diagonalizable over $\mathbb{C}$} and whose action on $(X,m_{X})$ is mixing. In this dynamical system we study the set of points $x\in X$ with a precompact orbit, written as $E(F,\infty)$, which is known to be a dense subset of $X$ of full Hausdorff dimension. We prove that $E(F,\infty)$ is indecomposable in the following sense: given any $y\in E(F,\infty)$, the set of $x\in E(F,\infty)$ for which $y\in\overline{F_{+}x}$, where $F_{+}$ denotes the positive ray in $F$, is uncountable and dense in $E(F,\infty)$. When the dimension of the neutral subgroup of $G$ with respect to $F$ is $1$ we demonstrate, for any $\varepsilon>0$, the existence of many points $x\in X$ whose orbit closure $\overline{F_{+}x}\subset X$ is compact and has Hausdorff dimension at least $\dim X-\varepsilon$.
@@ 1. Introduction
Take $y\in Y$ and consider $A(F,y):=A(F,\{y\})$ and $E(F,y):=E(F,\{y\})$. A natural question one could ask is how large the sets $A(F,y)$ and $E(F,y)$ can be. If the action admits an ergodic invariant probability measure, or has some hyperbolic behavior, some instances of this question can be answered. Namely, under the assumption of existence of an ergodic measure $\mu$ of full support the set $A(F,y)$ has full measure for any $y\in X\cup\{\infty\}$. Additionally, for many hyperbolic systems one can prove that the $\mu$-null sets $E(F,y)$ are thick, that is, their intersection with any nonempty open set has full Hausdorff dimension. For results of this kind see e.g. [D2] for toral endomorphisms, [U, Do] for Anosov flows, and [D1, KMa, Kl1, AGK] for partially hyperbolic flows on homogeneous space. Furthermore, in many cases these sets can be proved to be winning in the sense of Schmidt games (see e.g. [AGK, BFK, D1, D2, Ts1, Ts2, Wu1, Wu2]). Consequently in these cases it follows that both $A(F,y)$ and $E(F,y)$ are thick. Moreover, both full-measure and winning conditions are stable with respect to countable intersections; thus for any choice of countably many semigroups $F_{i}$ and points $y_{i}\in X\cup\{\infty\}$ for which the above conclusions can be established, it holds that both $\bigcap_{i}A(F_{i},y_{i})$ and $\bigcap_{i}E(F_{i},y_{i})$ are thick.
@@
Here and hereafter "dim" stands for Hausdorff dimension. In fact, even though it was not mentioned in [KR2], the proof actually demonstrates the thickness of the set in the left-hand side of (1.4). Also, by the standard "slicing" argument (see §2), it can be easily derived from Theorem 1.2 that for any $y\in E(F,\infty)$ the set $A(F_{+},y)\cap E(F_{+},\infty)$ is thick. The proof of the above theorem used the technique of continued fractions, and in fact the result can be interpreted within the framework of Diophantine approximation. Namely it constructs numbers that are badly approximable but not $\nu$-Dirichlet improvable, where $\nu$ is an irreducible norm on ${\mathbb{R}}^{2}$. See [KR1, KR2, KRS, AD] for more detail, and [MS] for some new results in that direction.
@@ 2. Reduction to the expanding subgroup
In this and the remaining part of the paper we let $G$, $\Gamma$ and $F$ be as in Theorem 1.3. Recall that $F=\{g_{t}=\exp(ta_{0}):t\in{\mathbb{R}}\}$, where ${a_{0}}\in{\mathfrak{g}}\smallsetminus\{0\}$ is $\operatorname{ad}$}-diagonalizable over $\mathbb{C}$. Define
$\begin{split}{\mathfrak{h}}&:=\left\{b\in{\mathfrak{g}}:\lim_{t\to-\infty}\operatorname{Ad}(g_{t})b=0\right\}\\ {\mathfrak{n}}&:=\left\{b\in{\mathfrak{g}}:\sup_{t\in{\mathbb{R}}}\|\operatorname{Ad}(g_{t})b\|<\infty\right\}\\ {\mathfrak{h}}^{-}&:=\left\{b\in{\mathfrak{g}}:\lim_{t\to\infty}\operatorname{Ad}(g_{t})b=0\right\}.\end{split}$
From the diagonalizability assumption, we have
${\mathfrak{g}}={\mathfrak{h}}\oplus{\mathfrak{n}}\oplus{\mathfrak{h}}^{-}.$ (2.1)
In fact, if one considers the subalgebras of $\mathfrak{g}$$\otimes_{\mathbb{R}}{\mathbb{C}}$ which are direct sums of eigenspaces for ${\operatorname{ad}}_{a_{0}}\otimes 1$ corresponding to eigenvalues with positive, zero and negative real part respectively, then one obtains the above algebra decomposition (2.1) simply by intersecting with $\mathfrak{g}$}.
Remark 2.1.
Note that $\mathfrak{n}$} is a proper subalgebra of $\mathfrak{g}$}, otherwise $\operatorname{Ad}(g_{t})$ would be a compact group and the mixing hypothesis would fail. Moreover, since $G$ admits a lattice, the modular function of $G$ is constant. Thus $\operatorname{{\operatorname{ad}}_{a_{0}}}$ has trace 0, and, therefore, both $\mathfrak{h}$} and $\mathfrak{h}$$^{-}$ are nonzero.
Definition 2.2.
[Expanding, contracting and neutral subgroups] We let $H$, $H^{-}$ and $Z$ be the connected Lie subgroups of $G$ associated to the algebras $\mathfrak{h}$}, $\mathfrak{h}$$^{-}$ and $\mathfrak{n}$} respectively. We fix $m_{H}$, a Haar measure on $H$. We further consider the connected subgroup $\tilde{H}$ associated to the Lie algebra ${\mathfrak{n}}\oplus{\mathfrak{h}}^{-}$ and a left Haar measure $m_{\tilde{H}}$ on $\tilde{H}$, scaled so that $m_{G}$ is a direct product of $m_{H}$ and $m_{\tilde{H}}$, see [B, Ch. VII, §9, Proposition 13] or [Kn, Theorem 8.32] for justification.
@@
The inner product on $\mathfrak{g}$} induces a norm which we denote by $\|\cdot\|$. Further, using the induced inner product on $\mathfrak{h}$}, $\mathfrak{h}$$^{-}$ and $\mathfrak{n}$} we obtain right-invariant Riemannian distance functions on $H$, $H^{-}$ and $Z$ denoted by $\operatorname{dist}_{H}$, $\operatorname{dist}_{H^{-}}$ and $\operatorname{dist}_{Z}$. In view of the commutation relation $\operatorname{Ad}\big(\exp(b)\big)=e^{{\operatorname{ad}}_{b}}$ for all $b\in{\mathfrak{g}}$, we have
$g_{t}\exp(b)g_{-t}=\exp(e^{t{\operatorname{ad}}_{a_{0}}}b)\ \ \text{for all $b\in{\mathfrak{g}}$ and }t\in{\mathbb{R}}.$ (2.2)
In particular, $H,H^{-}$ and $Z$ are invariant under conjugation by $g_{t}$.
For $g\in G$, $x\in X$, $a\in{\mathfrak{g}}$ and $r>0$, we use the notation
$B_{G}(g,r),\ B_{X}(x,r)\text{ and }B_{\mathfrak{g}}(a,r)$
to denote balls in $G$ with respect to $\operatorname{dist}_{G}$, balls in $X$ with respect to $\operatorname{dist}_{X}$ and balls in $\mathfrak{g}$} with respect to the norm induced by the inner product. When the base-point is the identity in $G$ or the zero element in $\mathfrak{g}$}, we use the notation $B_{G}(r)$ or $B_{\mathfrak{g}}(r)$.
@@
In view of (2.2), $\Phi_{t}$ restricts to a smooth group isomorphism of the groups $H$, $H^{-}$ and $Z$. We can thus define the Radon–Nikodym derivative
@@
Let $X=G/\Gamma$ and $F=\{g_{t}\in G:t\in{\mathbb{R}}\}$ be as in Theorem 1.3. Let $\sigma_{0}>0$, $x_{0}\in X$, and let $B$ be a countable set with ${FB}\subset X$ bounded; in particular, $B\subset E(F,\infty)$. Then the set of $h\in H$ with the properties
@@
is a diffeomorphism onto an open subset of $U$. It follows from the definition of $\mathfrak{n}$} that
[…]
Take $r_{0},C$ as in Lemma 2.4. By shrinking $r$ if needed we can assume that $r<r_{0}$. Apply Theorem 2.5 with $\sigma_{0}=r/C$ and $B=\{y\}$. This gives us a set
$S:=\left\{h\in B_{H}(r):hx_{0}\in E(F_{+},\infty)\cap A\big(F_{+},B_{Z}(\sigma_{0})y\big)\right\}$
@@
Lemma 2.4 thus shows that $g_{s}zh^{-}Sx_{0}\subset E(F_{+},\infty)$. Further, let $h\in S$, and suppose that $hx_{0}\in A(F_{+},zy)$ with $z\in B_{Z}({\sigma_{0}})$. This gives us an unbounded sequence $(t_{n})_{n\in{\mathbb{N}}}\subset{\mathbb{R}}_{\geq 0}$ with
@@
Choose a subsequence $(t_{n_{k}})$ and $t_{0}$ as in (2.4), and let $w=\Phi_{t_{0}}^{-1}(z^{-1})$. We then see that for any $h^{-}\in H^{-}$ and $s\in{\mathbb{R}}$
@@
Recall also that in this case we have ${\mathfrak{n}}={\mathbb{R}}a_{0}$ and $Z=F$. For any non-empty open $U\subset X$ we can choose a base point $x_{0}\in U$ and apply Theorem 2.5 with an arbitrary $\sigma_{0}>0$ and with $B\subset E(F,\infty)$ countable as above. This gives us a thick set of $h\in H$ such that, for every $y\in B$,
$hx_{0}\in E(F_{+},\infty)\cap A\big(F_{+},B_{Z}(\sigma_{0})y\big)=E(F_{+},\infty)\cap A(F_{+},y).$

### v2

@@ Abstract.
Let $X=G/\Gamma$ be a quotient of a real Lie group by a non-uniform lattice. Consider a one-parameter subgroup $F$ of $G$ that is $\operatorname{Ad}$-diagonalizable over $\mathbb{C}$} and whose action on $X$ is mixing. In this dynamical system we study the set of points $x\in X$ with a precompact orbit, written as $E(F,\infty)$, which is known to be a dense subset of $X$ of full Hausdorff dimension. We prove that $E(F,\infty)$ is indecomposable in the following sense: given any $y\in E(F,\infty)$, the set of $x\in E(F,\infty)$ for which $y\in\overline{F_{+}x}$, where $F_{+}$ denotes the positive ray in $F$, is uncountable and dense in $E(F,\infty)$. When the dimension of the neutral subgroup of $G$ with respect to $F$ is $1$ we demonstrate, for any $\varepsilon>0$, the existence of many points $x\in X$ whose orbit closure $\overline{F_{+}x}\subset X$ is compact and has Hausdorff dimension at least $\dim X-\varepsilon$.
@@ 1. Introduction
Take $y\in Y$ and consider $A(F,y):=A(F,\{y\})$ and $E(F,y):=E(F,\{y\})$. A natural question one could ask is how large the sets $A(F,y)$ and $E(F,y)$ can be. If the action admits an ergodic invariant probability measure, or has some hyperbolic behavior, some instances of this question can be answered. Namely, under the assumption of existence of an ergodic measure $\mu$ of full support the set $A(F,y)$ has full measure for any $y\in X\cup\{\infty\}$. Additionally, for many hyperbolic systems one can prove that the $\mu$-null sets $E(F,y)$ are thick, that is, their intersection with any nonempty open set has full Hausdorff dimension. For results of this kind see e.g. [D2] for toral endomorphisms, [U, Do] for Anosov flows, and [D1, KMa, Kl1, AGK] for partially hyperbolic flows on homogeneous space. Furthermore, in many cases these sets can be proved to be winning in the sense of Schmidt games [S]; see e.g. [AGK, BFK, D1, D2, Ts1, Ts2, Wu1, Wu2]). Consequently in these cases it follows that both $A(F,y)$ and $E(F,y)$ are thick. Moreover, both full-measure and winning conditions are stable with respect to countable intersections; thus for any choice of countably many semigroups $F_{i}$ and points $y_{i}\in X\cup\{\infty\}$ for which the above conclusions can be established, it holds that both $\bigcap_{i}A(F_{i},y_{i})$ and $\bigcap_{i}E(F_{i},y_{i})$ are thick.
@@
Here and hereafter "dim" stands for Hausdorff dimension. In fact, even though it was not mentioned in [KR2], the proof actually demonstrates the thickness of the set in the left-hand side of (1.4). Also, by the standard "slicing" argument (see §2), it can be easily derived from Theorem 1.2 that for any $y\in E(F,\infty)$ the set $A(F_{+},y)\cap E(F_{+},\infty)$ is thick. The proof of the above theorem used the technique of continued fractions, and in fact the result can be interpreted within the framework of Diophantine approximation. Namely it constructs numbers that are badly approximable but not $\nu$-Dirichlet improvable, where $\nu$ is an irreducible norm on ${\mathbb{R}}^{2}$ whose unit ball is not a parallelogram. See [AD, KR1, KR2, KRS] for more detail, and [MS] for some new results in that direction.
@@ 2. Reduction to the expanding subgroup
In this and the remaining part of the paper we let $G$, $\Gamma$ and $F$ be as in Theorem 1.3. Recall that $F=\{g_{t}=\exp(ta_{0}):t\in{\mathbb{R}}\}$, where ${a_{0}}\in{\mathfrak{g}}\smallsetminus\{0\}$ is $\operatorname{ad}$}-diagonalizable over $\mathbb{C}$.
Definition 2.1 (Expanding, contracting and neutral subgroups).
Consider
$\begin{split}{\mathfrak{h}}&:=\left\{b\in{\mathfrak{g}}:\lim_{t\to-\infty}\operatorname{Ad}(g_{t})b=0\right\},\\ {\mathfrak{h}}^{-}&:=\left\{b\in{\mathfrak{g}}:\lim_{t\to\infty}\operatorname{Ad}(g_{t})b=0\right\},\\ {\mathfrak{n}}&:=\left\{b\in{\mathfrak{g}}:\sup_{t\in{\mathbb{R}}}\|\operatorname{Ad}(g_{t})b\|<\infty\right\},\end{split}$
and let $H$, $H^{-}$ and $Z$ be the connected Lie subgroups of $G$ associated to the Lie algebras $\mathfrak{h}$}, $\mathfrak{h}$$^{-}$ and $\mathfrak{n}$} respectively. From the diagonalizability assumption, we have
${\mathfrak{g}}={\mathfrak{h}}\oplus{\mathfrak{n}}\oplus{\mathfrak{h}}^{-},$ (2.1)
hence $G$ is locally (in a small neighborhood of identity) a direct product of $H$, $H^{-}$ and $Z$. In fact, if one considers the subalgebras of $\mathfrak{g}$$\otimes_{\mathbb{R}}{\mathbb{C}}$ which are direct sums of eigenspaces for ${\operatorname{ad}}_{a_{0}}\otimes 1$ corresponding to eigenvalues with positive, zero and negative real part respectively, then one obtains the above Lie algebra decomposition (2.1) simply by intersecting with $\mathfrak{g}$}.
We remark that $\mathfrak{n}$} is a proper subalgebra of $\mathfrak{g}$}, otherwise $\operatorname{Ad}(g_{t})$ would belong to a compact group and the mixing hypothesis would fail. Moreover, since $G$ admits a lattice, the modular function of $G$ is constant. Thus ${\operatorname{ad}}_{a_{0}}$} has trace 0, and, therefore, both $\mathfrak{h}$} and $\mathfrak{h}$$^{-}$ (hence $H$ and $H^{-}$ as well) are non-trivial.
Definition 2.2 (Measures on $H$ and $\tilde{H}$).
We fix $m_{H}$, a Haar measure on $H$. We further consider the connected subgroup $\tilde{H}$ associated to the Lie algebra ${\mathfrak{n}}\oplus{\mathfrak{h}}^{-}$ and a left Haar measure $m_{\tilde{H}}$ on $\tilde{H}$, scaled so that $m_{G}$ is a direct product of $m_{H}$ and $m_{\tilde{H}}$, see [B, Ch. VII, §9, Proposition 13] or [Kn, Theorem 8.32] for justification.
@@
In view of the commutation relation $\operatorname{Ad}\big(\exp(b)\big)=e^{{\operatorname{ad}}_{b}}$ for all $b\in{\mathfrak{g}}$, we have
$\Phi_{t}\big(\exp(b)\big)=\exp(e^{t{\operatorname{ad}}_{a_{0}}}b)\ \ \text{for all $b\in{\mathfrak{g}}$ and }t\in{\mathbb{R}}.$
In particular, $H,H^{-}$ and $Z$ are invariant under $\Phi_{t}$, and the latter restricts to a smooth isomorphism of these groups. We can thus define the Radon–Nikodym derivative
@@
In what follows for a closed subgroup $P$ of $G$ we use the inner product on $\operatorname{Lie}(P)$ induced from $\mathfrak{g}$} to obtain the right-invariant Riemannian distance function $\operatorname{dist}_{P}$ on $P$. For $g\in P$, $x\in X$ and $r>0$ we use the notation
$B_{P}(g,r)\text{ and }B_{X}(x,r)$
to denote balls in $P$ with respect to $\operatorname{dist}_{P}$ and balls in $X$ with respect to $\operatorname{dist}_{X}$. When the base-point is the identity $e_{G}\in G$, we use the notation $B_{P}(r):=B_{P}(e_{G},r)$.
@@
Let $X=G/\Gamma$ and $F=\{g_{t}\in G:t\in{\mathbb{R}}\}$ be as in Theorem 1.3. Let $\rho>0$, $x_{0}\in X$, and let $B$ be a countable set with ${FB}\subset X$ bounded; in particular, $B\subset E(F,\infty)$. Then the set of $h\in H$ with the properties
@@
is a diffeomorphism onto an open subset of $U$. It follows from the definition of $\mathfrak{n}$} that there exists constants $\rho_{1},C>0$ such that for any $\rho<\rho_{1}$ and any unbounded sequence $(t_{n})_{n\in{\mathbb{N}}}\subset{\mathbb{R}}$ we have a subsequence $(t_{n_{k}})_{k\in{\mathbb{N}}}$ such that the sequence of maps $\Phi_{t_{k_{n}}}|_{B_{Z}(C\rho)}$ converges pointwise to some smooth immersion
$\Phi:B_{Z}(C\rho)\to Z$ (2.3)
whose image contains the ball $B_{Z}(\rho)$. Apply Theorem 2.5 with $\rho<\rho_{1}$. Further assume that $C\rho<r$. This gives us a set
$S:=\left\{h\in B_{H}(r):hx_{0}\in E(F_{+},\infty)\cap A\big(F_{+},B_{Z}(\rho)y\big)\right\}$
@@
Lemma 2.4 thus shows that $g_{s}zh^{-}Sx_{0}\subset E(F_{+},\infty)$. Further, let $h\in S$, and suppose that $hx_{0}\in A(F_{+},zy)$ with $z\in B_{Z}({\rho})$. This gives us an unbounded sequence $(t_{n})_{n\in{\mathbb{N}}}\subset{\mathbb{R}}_{\geq 0}$ with
@@
Choose a subsequence $(t_{n_{k}})$ and map $\Phi$ as in (2.3), and let $w=\Phi^{-1}(z^{-1})\in B_{Z}(C\rho)$. We then see that for any $h^{-}\in H^{-}$ and $s\in{\mathbb{R}}$
@@
Recall also that in this case we have ${\mathfrak{n}}={\mathbb{R}}a_{0}$ and $Z=F$. For any non-empty open $U\subset X$ we can choose a base point $x_{0}\in U$ and apply Theorem 2.5 with an arbitrary $\rho>0$ and with $B\subset E(F,\infty)$ countable as above. This gives us a thick set of $h\in H$ such that, for every $y\in B$,
$hx_{0}\in E(F_{+},\infty)\cap A\big(F_{+},B_{Z}(\rho)y\big)=E(F_{+},\infty)\cap A(F_{+},y).$